# SEMF: Supervised Expectation-Maximization Framework for Predicting Intervals

### v2

@@ Abstract
This work introduces the Supervised Expectation-Maximization Framework (SEMF), a versatile and model-agnostic framework that generates prediction intervals for datasets with complete or missing data. SEMF extends the Expectation-Maximization (EM) algorithm, traditionally used in unsupervised learning, to a supervised context, enabling it to extract latent representations for uncertainty estimation. The framework demonstrates robustness through extensive empirical evaluation across 11 tabular datasets, achieving—in some cases—narrower normalized prediction intervals and higher coverage than traditional quantile regression methods [^1] Code: https://github.com/Unco3892/semf-paper. Furthermore, SEMF integrates seamlessly with existing machine learning algorithms, such as gradient-boosted trees and neural networks, exemplifying its usefulness for real-world applications. The experimental results highlight SEMF’s potential to advance state-of-the-art techniques in uncertainty quantification.
@@ 1 Introduction
In the evolving field of machine learning (ML), the quest for models capable of predicting outcomes while quantifying the uncertainty of their predictions is critical. Estimating prediction uncertainty is particularly vital in high-stakes domains such as healthcare [1], finance [2], and autonomous systems [3], where prediction-based decisions can have significant consequences. Traditional approaches have primarily focused on point estimates, providing limited insight into the reliability of these predictions. This limitation underscores the necessity for frameworks that generate precise point predictions and robust prediction intervals. These intervals offer a range within which the true outcome is expected to lie with a specified probability, enabling a deeper understanding of prediction uncertainty. Consequently, research has increasingly focused on methodologies that extend beyond point estimation to include uncertainty quantification, thereby enhancing decision-making in applications that rely on predictive modeling [4].
In this paper, we introduce the Supervised Expectation-Maximization Framework (SEMF) based on the Expectation-Maximization (EM) algorithm [5]. Traditionally recognized as a clustering technique, EM is used for supervised learning in SEMF, allowing for both point estimates and prediction intervals using any ML model (model-agnostic). SEMF generates representations for latent or missing modalities, making it particularly relevant for incomplete multi-modal data. This paper details the methodology behind the framework and proposes a training algorithm based on Monte Carlo (MC) sampling, also used in variational inference for Variational Auto-Encoders (VAEs) [6, 7]. Comparisons with other models across several datasets showcase its contribution to the existing uncertainty quantification and predictive modeling literature [8].
@@ 2 Method
The EM algorithm has already been used for supervised learning tasks using specific models [10, 11, 12, 13], and, to the best of our knowledge, mainly focusing on point prediction through the use of Gaussian Mixture Models (GMMs). Our work differs by modifying the EM algorithm and MC sampling to generate prediction intervals with any ML model and, in theory, under any distribution. This results in SEMF, a framework consisting of an encoder-decoder architecture and a missing data simulator.
@@ 2.2.1 Example: $\mathcal{L}(\phi,\theta,\xi)$ under normality
When $x_{1,i}$ is missing, the sampling is enriched by simulated $x_{1,i,r}$. Eq. 16 selects $j^{*}$ from $I_{nm}$ based on $p_{\xi^{\prime}}(j|x_{2})$ according to Eq. 12 and Eq. 11. The above $D_{x}$ is a subset of the data for $\xi$ is only learned on the missing data part. After discussing the training steps in Subsection 2.3, the details are given in Algorithm 1 (Appendix A). We also note that the parameters above can be learned in parallel.
@@ 2.3 Training
For efficiency purpose, the training set, $\{1,\ldots,N\}$, is segmented into batches $\{b_{1},\ldots,b_{L}\}$ on which the index $i$ runs (and thus the denominator of Eq. 15 must be adapted accordingly). The process iterates for each batch until the maximum number of steps is reached or an early stopping criterion is satisfied. The framework requires tuning hyper-parameters such as the number of MC samples $R$, the number of latent nodes $m_{k}$, and the standard deviation $\sigma_{k}$ of $Z_{k}$. Monitoring the point prediction on a hold-out validation is important to combat overfitting and terminate the training early with a patience hyper-parameter. Moreover, due to the generative nature of SEMF, the variation resulting from the initial random seed is measured in Subsection 4.2. Additionally. the model-specific hyper-parameters ($p_{\phi}$, $p_{\theta}$ and $p_{\xi}$) are also discussed in the same Subsection.
@@ 3.1 Latent Representation Learning
Latent representation learning involves modeling hidden variables from observed data for various ML tasks, most notably Auto-Encoders (AEs) and VAEs. AEs are neural networks reconstructing inputs by learning an intermediate latent representation. The leftmost part of AE is the encoder $g_{\phi}(\cdot)$, which embeds the input $x$ into a latent variable $z$. It is passed to the rightmost part of AE, the decoder $f_{\theta}(\cdot)$, for the reconstruction $\hat{x}=f_{\theta}(z)$ of $x$ through $z$. The parameters $\phi$ and $\theta$ are trained by minimizing the reconstruction loss. Unlike AEs, which focus on reconstruction, VAEs use variational methods to fit distributions of latent variables and of the output [7]. Given a sample $x$ with a latent variable of $z$, VAEs learn the marginal likelihood, $p_{\theta}(x)$. Direct maximization of this likelihood is difficult [14]. Thus, alternative methods exist, such as maximizing the evidence lower bound (ELBO), which provides guarantees on the log-likelihood [15].
For supervised and semi-supervised tasks, latent representation learning can include task-specific predictions [16]. More specifically, models such as AEs follow the classical encoder-decoder objective while training a predictor $h_{\psi}(z)$ through an additional layer or model to estimate the output $y$. This dual objective helps in learning more task-relevant embeddings [17, 18]. Semi-supervised VAEs are similar, with the distinction that they couple the reconstruction loss of the unlabeled data with a variational approximation of latent variables. This is effective even with sparse labels [19, 20]. Similarly, the EM algorithm adapts well to minimal supervision [21] and using labeled and unlabeled data in semi-supervised settings for both single and multiple modalities [12, 22].
@@ 3.2 Prediction Intervals
Crucial for estimating uncertainty, prediction intervals in regression are often derived using methods such as Bayesian approaches [23, 24, 25], ensemble techniques [26, 27], or quantile regression [28, 29]. Additionally, these methods can be complemented with conformal prediction, a framework for calibrating any point predictor to produce prediction intervals [30, 31], making it highly relevant for enhancing reliability in applications requiring rigorous uncertainty quantification.
@@ 3.3 Missing Data
Managing missing values is a pivotal aspect when dealing with real-world data. Naive methods such as discarding instances or mean/median-imputation may be infeasible or carry the risk of changing the data distribution [35, 36] introducing bias if the missing values are not missing at random (MAR) or completely at random (MCAR). The chosen technique for handling missing data should adhere to the dataset’s characteristics and mechanisms behind the missing data [37].
More advanced approaches, like the Iterative Imputer from scikit-learn [38], an implementation of Multiple Imputation by Chained Equations (MICE) [39], expand on the simple imputations by iteratively modeling each feature with missing values as a function of other features [40, 41]. Due to its complexity, MICE is best compared with alternatives such as K-means clustering and artificial neural networks (ANNs). K-means assigns missing values based on cluster centroids [42] as opposed to ANNs, which learn complex, non-linear relationships between variables [43]. ANNs effectively predict or reconstruct missing values and thus are particularly useful in datasets where relationships between variables are intricate and not easily captured by straightforward imputation methods. Accordingly, ANNs have demonstrated superior performance over MICE and GMM (with EM for missing values) in scenarios where a large proportion of the data is missing [44].
@@ 4.1 Datasets
We systematically curate a subset of datasets from the OpenML-CTR23 [45] benchmark suite to evaluate and carry out our experiments. Initially comprising 35 datasets, we apply an exclusion criteria to refine this collection to 11 datasets. The details and overview are in Appendix B. We remove duplicated rows from all the datasets and carry out the standardization (scaling) of all predictors, including the outcome, which we transformed to have zero means and unit variances. The features of these datasets are then treated as separate inputs to SEMF. In the second stage of our experiments, we artificially introduce 50% missing values in our datasets for any predictor except for the first feature of a randomly chosen row, which emulates MAR or MCAR data. In all our datasets, 70% of the data is used to train all models, 15% as a hold-out validation set to monitor SEMF’s performance, and 15% to evaluate the models. To combat overfitting, baseline models that benefit from early stopping are allocated another 15% from the training data. Lastly, it is essential to note that all data in SEMF are processed batch-wise, without employing mini-batch training, to ensure consistency and stability in the training process.
@@ 4.2 Models
Our baseline consists of both point and quantile regression eXtreme Gradient Boosting (XGBoost) [46], Extremely Randomized Trees (ET) [47], and neural networks [48], depicted in Table 1. To ensure consistency in our experimental setup, we align the families and hyper-parameters of $p_{\phi}$ and $p_{\theta}$ with our baseline models. For example, in the case of XGBoost in SEMF, we use $K$ XGBoost’s $g_{\phi_{k}}(x_{k})$, one for each input $x_{k}$, $k=1,\ldots,K$, and one XGBoost for $f_{\theta}(z)$ with the same hyper-parameters. We refer to the SEMF’s adoption of these models as MultiXGBs, MultiETs, and MultiMLPs. When establishing prediction intervals, we set our uncertainty tolerance at 5% for both the baseline and SEMF (Eq. 18). The missing data simulator, $p_{\xi}$, is constructed using a shallow neural network, which employs the $\mathrm{SELU}$ activation function [49]. It is experimented with two distinct node counts: 50 and 100.
@@
Trees: 100, Maximum depth: 5, Early stopping steps: 10
@@ 4.3 Metrics
The evaluation of point predictions employs Root Mean Squared Error (RMSE), Mean Absolute Error (MAE), and R-squared ($\text{R}^{2}$). For prediction intervals, the chosen metrics, following [52, 53], are the Prediction Interval Coverage Probability (PICP), Mean Prediction Interval Width (MPIW), and Normalized Mean Prediction Interval Width (NMPIW) as described in Subsection D.1. This paper also evaluates SEMF on our new metric, termed Coverage-Width Ratio (CWR),
@@ 5 Results
We trained and tested 330 models corresponding to the three model types—MultiXGBs, MultiETs, and MultiMLPs—across 11 datasets with both complete and 50% missing data, with five seeds for each combination. Table 2 and Table 3 provide the mean and standard deviation for our metrics aggregated over the five seeds. All the results are in Appendix E. Below, we discuss the performances with complete data and with incomplete data.
@@ 5.1 Complete Data
The results of models with complete data are in Table 2. Overall, MultiXGBs and MultiMLPs performed generally well in producing intervals compared to their baselines as shown by positive $\Delta$CWR and $\Delta$PICP, notably, for instance, on cpu_activity, naval_propulsion_plant, and energy_efficiency. Interestingly, MultiMLPs also attained good relative performance improvements for point prediction, which can indicate that the chosen $\sigma_{k}$ was too low for generating performant prediction intervals. MultiETs attained mixed results as, despite significant improvements in $\Delta$CWR, it often failed to beat the baseline on $\Delta$PICP.
Table 2: Test results for all models with complete data at 95% quantiles aggregated over five seeds. For each metric, the mean and standard deviation of the performance across the seeds are separated by ±. Performance over the baseline is highlighted in bold.
@@
space_ga -1%±6% 6%±2% -7%±6% 0.89±0.01 1.80±0.12 0.20±0.01 -9%±1% -10%±2% 0.60±0.01
cpu_activity 25%±7% 9%±1% 12%±5% 0.89±0.01 0.37±0.01 0.07±0.00 21%±1% -1%±3% 0.98±0.00
naval_propulsion_plant 156%±15% 8%±1% 58%±2% 0.96±0.00 0.39±0.01 $0$14%±4% -12%±2% 0.99±0.00
miami_housing 61%±6% -7%±2% 43%±3% 0.82±0.01 0.45±0.01 0.06±0.00 4%±5% 3%±3% 0.91±0.01
kin8nm 10%±3% 5%±2% 4%±4% 0.90±0.01 1.94±0.07 0.38±0.01 -20%±2% -22%±2% 0.63±0.01
concrete_compressive_strength 13%±5% 26%±6% -11%±8% 0.91±0.01 1.25±0.03 0.25±0.01 -26%±7% -40%±13% 0.85±0.01
cars 18%±9% 15%±4% 2%±9% 0.89±0.02 0.70±0.05 0.12±0.01 -3%±3% -1%±4% 0.95±0.00
energy_efficiency 247%±99% 22%±7% 60%±20% 0.88±0.02 0.16±0.04 0.05±0.01 -15%±25% -16%±19% 1.00±0.00
california_housing 31%±5% 0%±1% 23%±4% 0.88±0.01 1.17±0.02 $0$1%±1% -1%±1% 0.81±0.00
airfoil_self_noise 61%±45% 4%±10% 31%±18% 0.81±0.07 0.89±0.22 0.19±0.05 -64%±36% -73%±46% 0.86±0.05
QSAR_fish_toxicity 10%±4% 11%±4% -1%±3% 0.76±0.01 1.46±0.03 0.23±0.00 3%±4% 1%±5% 0.55±0.02
@@
space_ga 9%±4% -2%±1% 9%±4% 0.91±0.01 2.04±0.07 0.22±0.01 -16%±2% -18%±3% 0.54±0.02
cpu_activity 27%±3% -8%±0% 27%±2% 0.90±0.01 0.47±0.01 $0$14%±2% -20%±2% 0.98±0.00
naval_propulsion_plant 219%±36% -9%±0% 71%±3% 0.91±0.00 0.66±0.07 0.19±0.02 -320%±45% -409%±40% 0.96±0.01
miami_housing 69%±4% -13%±0% 48%±1% 0.86±0.00 0.65±0.02 $0$10%±2% -20%±2% 0.90±0.00
kin8nm 10%±2% -9%±0% 17%±2% 0.88±0.01 2.27±0.05 0.44±0.01 -35%±2% -38%±2% 0.48±0.01
concrete_compressive_strength 9%±10% -12%±2% 19%±9% 0.81±0.02 1.18±0.11 0.24±0.02 -67%±7% -94%±10% 0.76±0.02
cars 11%±7% 16%±2% -5%±7% 0.83±0.02 0.53±0.04 0.09±0.01 3%±4% 1%±2% 0.95±0.00
energy_efficiency -4%±6% 0%±8% -5%±15% 0.65±0.04 0.08±0.01 0.02±0.00 3%±4% 0%±2% 1.00±0.00
california_housing 31%±2% -7%±0% 29%±1% 0.90±0.00 1.70±0.03 0.41±0.01 -15%±1% -21%±1% 0.71±0.01
airfoil_self_noise 18%±28% -15%±5% 25%±18% 0.83±0.05 1.22±0.31 0.26±0.07 -117%±50% -141%±57% 0.80±0.09
QSAR_fish_toxicity 12%±2% -9%±1% 19%±2% 0.82±0.01 1.81±0.02 $0$6%±1% -11%±1% 0.53±0.01
@@
space_ga 8%±1% 0%±3% 7%±3% 0.81±0.02 1.25±0.02 0.14±0.00 0%±1% -1%±1% 0.75±0.01
cpu_activity 8%±3% -9%±4% 15%±6% 0.72±0.02 0.24±0.01 0.05±0.00 7%±2% 5%±2% 0.98±0.00
naval_propulsion_plant 21%±20% 0%±3% 16%±15% 0.92±0.01 0.22±0.03 0.07±0.01 -45%±28% -36%±23% 1.00±0.00
miami_housing 9%±4% 2%±3% 6%±6% 0.81±0.02 0.41±0.03 $0$3%±2% 4%±2% 0.91±0.00
kin8nm 2%±8% 8%±6% -6%±12% 0.81±0.02 0.68±0.03 0.13±0.01 7%±3% 7%±2% 0.93±0.00
concrete_compressive_strength 98%±15% -23%±8% 61%±5% 0.56±0.04 0.28±0.01 0.06±0.00 12%±6% 15%±4% 0.91±0.01
cars 0%±15% 6%±13% -9%±33% 0.76±0.09 0.40±0.10 0.07±0.02 -3%±3% 0%±3% 0.95±0.00
energy_efficiency 82%±20% 25%±26% 31%±12% 0.62±0.05 0.06±0.00 0.02±0.00 32%±4% 33%±3% 1.00±0.00
california_housing -14%±3% 8%±1% -26%±4% 0.89±0.01 1.31±0.02 0.31±0.01 7%±0% 9%±1% 0.82±0.00
airfoil_self_noise 74%±46% 0%±7% 39%±17% 0.75±0.02 0.35±0.01 0.08±0.00 18%±4% 16%±5% 0.97±0.00
QSAR_fish_toxicity -6%±4% 11%±4% -19%±5% 0.76±0.04 1.49±0.06 0.23±0.01 4%±5% 7%±5% 0.55±0.04

### v3

@@ Abstract
This work introduces the Supervised Expectation-Maximization Framework (SEMF), a versatile and model-agnostic approach for generating prediction intervals in datasets with complete or missing data. SEMF extends the Expectation-Maximization algorithm, traditionally used in unsupervised learning, to a supervised context, leveraging latent variable modeling for uncertainty estimation. Extensive empirical evaluations across 11 tabular datasets show that SEMF often achieves narrower normalized prediction intervals and higher coverage rates than traditional quantile regression methods [^1] Code: https://github.com/Unco3892/semf-paper. Furthermore, SEMF can be integrated with machine learning models like gradient-boosted trees and neural networks, highlighting its practical applicability. The results indicate that SEMF enhances uncertainty quantification, particularly in scenarios with complete data.
@@ 1 Introduction
In the evolving field of machine learning (ML), the quest for models able to predict outcomes while quantifying the uncertainty of their predictions is critical. The ability to estimate prediction uncertainty is particularly vital in high-stakes domains such as healthcare [1], finance [2], and autonomous systems [3], where prediction-based decisions have important consequences. Traditional approaches have primarily focused on point estimates, with little to no insight into prediction reliability. This limitation underscores the need for frameworks that can generate both precise point predictions and robust prediction intervals. Such intervals provide a range within which the true outcome is expected to lie with a fixed probability, offering a finer understanding of prediction uncertainty. This need has spurred research into methodologies that extend beyond point estimation to include uncertainty quantification, thereby enabling more informed decision-making in applications reliant on predictive modeling [4].
In this paper, we introduce the Supervised Expectation-Maximization Framework (SEMF) based on the Expectation-Maximization (EM) algorithm [5]. Traditionally recognized as a clustering technique, EM is used for supervised learning in SEMF, allowing for both point estimates and prediction intervals using any ML model (model-agnostic). SEMF generates representations for latent or missing modalities, which can be relevant for incomplete data and holds potential for multi-modal data applications, though multi-modal settings are left for future exploration. This paper details the methodology behind the framework and proposes a training algorithm based on Monte Carlo (MC) sampling, also used in variational inference for Variational Auto-Encoders (VAEs) [6, 7]. SEMF differs from prominent supervised EM approaches such as [8], which focus on point prediction using Gaussian Mixture Models (GMMs). Additionally, our method operates in a frequentist paradigm, directly maximizing the likelihood function through iterative EM steps without integrating over posterior distributions. Although SEMF can be extended to a Bayesian framework as its likelihood component, this extension lies beyond the scope of this paper.
@@ 2.2.1 Example: $\mathcal{L}(\phi,\theta,\xi)$ under normality
When $x_{1,i}$ is missing, the sampling is enriched by simulated $x_{1,i,r}$. Eq. 16 selects $j^{*}$ from $I_{nm}$ based on $p_{\xi^{\prime}}(j|x_{2})$ according to Eq. 12 and Eq. 11. The $D_{x}$ above is a subset of the data for $\xi$, learned on the missing data part. We also note that learning the parameters above is parallelizable.
@@ 2.3 Training
For efficiency purpose, the training set, $\{1,\ldots,N\}$, is segmented into batches $\{b_{1},\ldots,b_{L}\}$ on which the index $i$ runs (and thus the denominator of Eq. 15 must be adapted accordingly). The process iterates for each batch until the maximum number of steps is reached or an early stopping criterion is satisfied. The full details are given in Algorithm 1 (Appendix A). The framework requires tuning hyper-parameters such as the number of MC samples $R$, the number of latent nodes $m_{k}$, and the standard deviation $\sigma_{k}$ of $Z_{k}$. Monitoring the point prediction on a hold-out validation is important to combat overfitting and terminate the training early with a patience hyper-parameter. Moreover, due to the generative nature of SEMF, the variation resulting from the initial random seed is measured in Subsection 4.2. Additionally. the model-specific hyper-parameters ($p_{\phi}$, $p_{\theta}$ and $p_{\xi}$) are also discussed in the same Subsection.
@@ 3.1 Latent Representation Learning
Latent representation learning involves modeling hidden variables from observed data for various ML tasks, most notably Auto-Encoders (AEs) and VAEs. AEs are neural networks that reconstruct inputs by learning an intermediate latent representation. The encoder $g_{\phi}(\cdot)$ in an AE embeds the input $x$ into a latent variable $z$, which then passes through the decoder $f_{\theta}(\cdot)$, reconstructing $x$ as $\hat{x}=f_{\theta}(z)$. The training process optimizes the parameters $\phi$ and $\theta$ by minimizing the reconstruction loss. Unlike AEs, which focus on reconstruction, VAEs use variational methods to fit distributions of latent variables and the output [7]. Given a sample $x$ with a latent variable $z$, VAEs model the marginal likelihood $p_{\theta}(x)$. However, directly maximizing this likelihood is difficult [10]. Thus, alternative methods exist, such as maximizing the evidence lower bound (ELBO), which provides guarantees on the log-likelihood [11].
For supervised and semi-supervised tasks, latent representation learning can include task-specific predictions [12]. More specifically, models such as AEs follow the classical encoder-decoder objective while training a predictor $h_{\psi}(z)$ through an additional layer or model to estimate the output $y$. This dual objective helps in learning more task-relevant embeddings [13, 14]. Semi-supervised VAEs are similar, with the distinction that they couple the reconstruction loss of the unlabeled data with a variational approximation of latent variables. This is effective even with sparse labels [15, 16].
The EM algorithm has already been used for supervised learning tasks using specific models [8, 17, 18], where the goal has been point prediction with GMMs. Similarly, the EM algorithm adapts well to minimal supervision [19] and using labeled and unlabeled data in semi-supervised settings for both single and multiple modalities [20, 21]. Our work differs by modifying using MC sampling to generate prediction intervals with any ML model and, in theory, under any distribution.
@@ 3.2 Prediction Intervals
Crucial for estimating uncertainty, prediction intervals in regression are often derived using methods such as Bayesian approaches [22, 23, 24], ensemble techniques [25, 26, 27], or quantile regression [28, 29]. Additionally, these methods can be complemented with conformal prediction, a framework for calibrating any point predictor to produce prediction intervals [30, 31], making it highly relevant for enhancing reliability in applications requiring rigorous uncertainty quantification.
@@ 3.3 Missing Data
Managing missing values is a pivotal aspect when dealing with real-world data. Naive methods such as discarding instances or mean/median-imputation may be infeasible or carry the risk of changing the data distribution [35, 36]. The chosen technique for handling missing data should adhere to the dataset’s characteristics and mechanisms behind the missing data [37]. More advanced approaches, like the Iterative Imputer from scikit-learn [38], an implementation of Multiple Imputation by Chained Equations (MICE) [39], expand on the simple imputations by iteratively modeling each feature with missing values as a function of other features [40, 41]. Due to its complexity, MICE is best compared with alternatives such as K-means clustering and artificial neural networks (ANNs). K-means assigns missing values based on cluster centroids [42] as opposed to ANNs, which learn complex, non-linear relationships between variables [43]. ANNs effectively predict or reconstruct missing values and thus are particularly useful in datasets where relationships between variables are intricate and not easily captured by straightforward imputation methods. Accordingly, ANNs have demonstrated superior performance over MICE and GMM (with EM for missing values) in scenarios where a large proportion of the data is missing [44].
@@ 4.1 Datasets
We systematically curate a subset of datasets from the OpenML-CTR23 [45] benchmark suite to evaluate and carry out our experiments. Initially comprising 35 datasets, we apply an exclusion criteria to refine this collection to 11 datasets. The details and overview are in Appendix B. We remove duplicated rows from all the datasets and carry out the standardization (scaling) of all predictors, including the outcome, which we transformed to have zero means and unit variances. The features of these datasets are then treated as separate inputs to SEMF. In the second stage of our experiments, we artificially introduce 50% missing values in our datasets for any predictor except for the first feature of a randomly chosen row, which emulates missing completely at random (MCAR) data. In all our datasets, 70% of the data is used to train all models, 15% as a hold-out validation set to monitor SEMF’s performance, and 15% to evaluate the models. To combat overfitting, baseline models that benefit from early stopping are allocated another 15% from the training data. Lastly, it is essential to note that all data in SEMF are processed batch-wise, without employing mini-batch training, to ensure consistency and stability in the training process.
@@ 4.2 Models
Our baseline consists of both point and quantile regression eXtreme Gradient Boosting (XGBoost) [46], Extremely Randomized Trees (ET) [47], and neural networks [48], all summarized and depicted in Table 1. To ensure consistency in our experimental setup, we align the families and hyper-parameters of $p_{\phi}$ and $p_{\theta}$ with our baseline models. For example, in the case of XGBoost in SEMF, we use $K$ XGBoosts, $g_{\phi_{k}}(x_{k})$, one for each input $x_{k}$, $k=1,\ldots,K$, and one XGBoost for $f_{\theta}(z)$ with the same hyper-parameters. We refer to the SEMF’s adoption of these models as MultiXGBs, MultiETs, and MultiMLPs. When establishing prediction intervals, we conformalize our prediction intervals according to [34] at an uncertainty tolerance of 5% for both the baseline and SEMF (Eq. 18). The missing data simulator, $p_{\xi}$, is constructed using a shallow neural network, which employs the $\mathrm{SELU}$ activation function [49]. It is experimented with two distinct node counts: 50 and 100.
@@
Trees: 100, Maximum depth: 6, Early stopping steps: 10
@@ 4.3 Metrics
The evaluation of point predictions employs Root Mean Squared Error (RMSE), Mean Absolute Error (MAE), and R-squared ($\text{R}^{2}$). For prediction intervals, the chosen metrics, following [52] and [53], are the Prediction Interval Coverage Probability (PICP), and Normalized Mean Prediction Interval Width (NMPIW) as described in Subsection D.1. This paper also evaluates SEMF on our new metric, termed Coverage-Width Ratio (CWR),
@@ 5 Results
We trained and tested 330 models corresponding to the three model types—MultiXGBs, MultiETs, and MultiMLPs—across 11 datasets with both complete and 50% missing data, using five seeds for each combination. Table 2 and Table 3 present the mean and standard deviation for our metrics aggregated over the five seeds. Appendix E includes the results from each individual run. For comparison, we also present the non-conformalized prediction intervals in Appendix F, though we solely discuss the results for conformalized intervals on both complete and incomplete data.
@@ 5.1 Complete Data
The results of models with complete data are in Table 2. Overall, MultiXGBs and MultiMLPs performed generally well in producing intervals compared to their baselines, as shown by positive $\Delta$CWR and $\Delta$NMPIW while achieving similar $\Delta$PICP to the baselines. Notably, all models attained suitable intervals on naval_propulsion_plant and energy_efficiency. Interestingly, MultiMLPs also attained good relative performance improvements for point prediction, which can indicate that the chosen $\sigma_{k}$ was too low for generating performant prediction intervals. MultiETs attained mixed results, with significant improvements in $\Delta$CWR for other datasets such as naval_propulsion_plant, but its overall performance remains less conclusive.
Table 2: Test results for all models with complete data at 95% quantiles aggregated over five seeds. For each metric, the mean and standard deviation of the performance across the seeds are separated by $\pm$. Performance over the baseline is highlighted in bold.
@@
Dataset $\Delta$CWR $\Delta$PICP $\Delta$NMPIW PICP NMPIW $\Delta$RMSE $\Delta$MAE $\text{R}^{2}$
@@
space_ga 6%$\pm$3-1%$\pm$07%$\pm$20.94$\pm$0.01 0.26$\pm$0.01 -9%$\pm$1-10%$\pm$20.60$\pm$0.01
cpu_activity 16%$\pm$61%$\pm$112%$\pm$50.94$\pm$0.01 0.09$\pm$0.00 21%$\pm$1-1%$\pm$20.98$\pm$0.00
naval_propulsion_plant 172%$\pm$140%$\pm$163%$\pm$20.95$\pm$0.01 $0$14%$\pm$4-12%$\pm$20.99$\pm$0.00
miami_housing -7%$\pm$30%$\pm$0-8%$\pm$30.95$\pm$0.00 0.13$\pm$0.00 4%$\pm$43%$\pm$30.91$\pm$0.01
kin8nm 6%$\pm$21%$\pm$05%$\pm$20.94$\pm$0.01 0.45$\pm$0.01 -20%$\pm$1-22%$\pm$10.63$\pm$0.01
concrete_compressive_strength 41%$\pm$12-3%$\pm$231%$\pm$70.94$\pm$0.02 0.31$\pm$0.01 -26%$\pm$6-40%$\pm$110.85$\pm$0.01
cars 40%$\pm$18-3%$\pm$130%$\pm$80.91$\pm$0.01 0.13$\pm$0.01 -3%$\pm$3-1%$\pm$30.95$\pm$0.00
energy_efficiency 222%$\pm$45-4%$\pm$370%$\pm$30.92$\pm$0.02 0.05$\pm$0.01 -15%$\pm$22-16%$\pm$171.00$\pm$0.00
california_housing -1%$\pm$30%$\pm$0-2%$\pm$40.95$\pm$0.00 0.42$\pm$0.01 -1%$\pm$1-1%$\pm$10.81$\pm$0.00
airfoil_self_noise 21%$\pm$21-1%$\pm$215%$\pm$150.97$\pm$0.02 0.36$\pm$0.06 -64%$\pm$32-73%$\pm$410.86$\pm$0.05
QSAR_fish_toxicity 34%$\pm$8-5%$\pm$329%$\pm$50.87$\pm$0.02 0.33$\pm$0.01 3%$\pm$31%$\pm$40.55$\pm$0.02
@@
space_ga -6%$\pm$31%$\pm$1-7%$\pm$30.96$\pm$0.00 0.29$\pm$0.01 -15%$\pm$2-18%$\pm$30.54$\pm$0.02
cpu_activity 11%$\pm$2-4%$\pm$013%$\pm$20.94$\pm$0.00 $0$14%$\pm$1-20%$\pm$20.98$\pm$0.00
naval_propulsion_plant 137%$\pm$221%$\pm$057%$\pm$40.96$\pm$0.00 0.27$\pm$0.02 -316%$\pm$45-406%$\pm$430.96$\pm$0.01
miami_housing -9%$\pm$1-1%$\pm$0-9%$\pm$10.95$\pm$0.00 $0$10%$\pm$2-20%$\pm$20.90$\pm$0.00
kin8nm -10%$\pm$11%$\pm$1-12%$\pm$20.94$\pm$0.01 0.53$\pm$0.01 -34%$\pm$2-38%$\pm$20.48$\pm$0.01
concrete_compressive_strength -8%$\pm$5-3%$\pm$2-6%$\pm$60.89$\pm$0.02 0.31$\pm$0.02 -67%$\pm$6-94%$\pm$80.76$\pm$0.01
cars -25%$\pm$50%$\pm$3-34%$\pm$140.92$\pm$0.02 0.15$\pm$0.01 3%$\pm$41%$\pm$20.95$\pm$0.00
energy_efficiency 10%$\pm$6-4%$\pm$212%$\pm$60.94$\pm$0.02 0.06$\pm$0.00 3%$\pm$21%$\pm$11.00$\pm$0.00
california_housing 1%$\pm$2-2%$\pm$03%$\pm$20.95$\pm$0.00 0.55$\pm$0.01 -15%$\pm$1-21%$\pm$10.71$\pm$0.01
airfoil_self_noise -16%$\pm$10-3%$\pm$2-17%$\pm$120.96$\pm$0.02 0.43$\pm$0.03 -118%$\pm$43-141%$\pm$490.80$\pm$0.08
QSAR_fish_toxicity -6%$\pm$9-5%$\pm$2-2%$\pm$120.88$\pm$0.02 0.36$\pm$0.02 -6%$\pm$1-11%$\pm$10.53$\pm$0.01
@@
space_ga 6%$\pm$3-1%$\pm$16%$\pm$20.95$\pm$0.01 0.23$\pm$0.00 0%$\pm$1-1%$\pm$10.75$\pm$0.01
cpu_activity -7%$\pm$60%$\pm$1-9%$\pm$70.95$\pm$0.00 0.10$\pm$0.00 7%$\pm$25%$\pm$20.98$\pm$0.00
naval_propulsion_plant 4%$\pm$161%$\pm$00%$\pm$140.96$\pm$0.00 0.08$\pm$0.01 -45%$\pm$25-36%$\pm$201.00$\pm$0.00
miami_housing -38%$\pm$30%$\pm$1-61%$\pm$90.95$\pm$0.00 0.15$\pm$0.01 -3%$\pm$24%$\pm$20.91$\pm$0.00
kin8nm 8%$\pm$50%$\pm$18%$\pm$50.95$\pm$0.01 0.20$\pm$0.01 7%$\pm$27%$\pm$20.93$\pm$0.00
concrete_compressive_strength 11%$\pm$6-2%$\pm$211%$\pm$60.94$\pm$0.02 0.29$\pm$0.03 12%$\pm$515%$\pm$40.91$\pm$0.01
cars 3%$\pm$12-3%$\pm$34%$\pm$110.92$\pm$0.02 0.13$\pm$0.01 -3%$\pm$30%$\pm$30.95$\pm$0.00
energy_efficiency 53%$\pm$210%$\pm$434%$\pm$100.96$\pm$0.02 0.05$\pm$0.00 32%$\pm$333%$\pm$31.00$\pm$0.00
california_housing -12%$\pm$30%$\pm$1-14%$\pm$40.95$\pm$0.00 0.43$\pm$0.01 7%$\pm$09%$\pm$10.82$\pm$0.00
airfoil_self_noise 68%$\pm$28-2%$\pm$140%$\pm$110.97$\pm$0.01 0.18$\pm$0.01 18%$\pm$416%$\pm$50.97$\pm$0.00
QSAR_fish_toxicity 5%$\pm$101%$\pm$13%$\pm$100.89$\pm$0.03 0.35$\pm$0.04 4%$\pm$57%$\pm$40.55$\pm$0.04